# Configuring Agentic AI Coding Tools: An Exploratory Study

### v1

@@ 8. Conclusions
Our study is the first to provide a comprehensive overview of the configuration mechanisms used by agentic AI coding tools in software engineering. We systematically identified eight configuration mechanisms across five agentic AI coding tools and analyzed their adoption in 2,926 GitHub repositories. Three findings stand out. First, Context Files dominate and are often the only mechanism present, with AGENTS.md emerging as an interoperable standard. Second, advanced mechanisms such as Skills and Subagents remain shallowly adopted, with Skills used predominantly as static documents rather than executable workflows. Third, distinct configuration cultures are forming around different tools. For practitioners, these findings suggest that AGENTS.md is the natural starting point for configuring agentic tools, especially in multi-tool environments, and that Skills offer untapped potential for encoding recurring workflows beyond static instructions. Tool providers should consider improving onboarding and documentation for advanced mechanisms, given the gap between their expressive power and current adoption. For researchers, controlled studies are needed to determine whether richer configuration, such as Skills with executable resources, yields measurable improvements over Context Files alone, and how conflicts between configuration artifacts in multi-tool repositories can be detected and resolved. Longitudinal research needs to track how configuration patterns evolve.

### v2

@@ 8. Conclusions
Our study is the first to provide a comprehensive overview of the configuration mechanisms used by agentic AI coding tools in software engineering. We systematically identified eight configuration mechanisms across five agentic AI coding tools and analyzed their adoption in 2,923 GitHub repositories. Three findings stand out. First, Context Files dominate and are often the only mechanism present, with AGENTS.md emerging as an interoperable standard. Second, advanced mechanisms such as Skills and Subagents remain shallowly adopted, with Skills used predominantly as static documents rather than executable workflows. Third, distinct configuration cultures are forming around different tools. For practitioners, these findings suggest that AGENTS.md is the natural starting point for configuring agentic tools, especially in multi-tool environments, and that Skills offer untapped potential for encoding recurring workflows beyond static instructions. Tool providers should consider improving onboarding and documentation for advanced mechanisms, given the gap between their expressive power and current adoption. For researchers, controlled studies should determine whether richer configuration yields measurable improvements over Context Files alone, and longitudinal research should track how configuration patterns evolve.